OpenStudy (collegemanatopenstudy):

The side of an equilateral triangle is 2 inches shorter than the side of a square. The perimeter of the square is 30 inches more than the perimeter of the triangle. Find the length of a side of the triangle. Hows the answer 22Inch???

OpenStudy (pythagoras123):



OpenStudy (pythagoras123):

\[4x = 30+3(x-2)\] \[4x = 30 + 3x -6\] \[x=24\]

OpenStudy (mathmale):

Why not experiment a bit: substitute x=22, and then x=24, into 4x=30+3(x-2). Does this equation turn out to be true in either case?